OpenStudy (anonymous):

Divide. Assume the denominator does not equal zero. (18x^5y^3+6x^2y^4+24x^6y) ÷ 3xy Can you please show me how to do this with work?

OpenStudy (anonymous):

\[\frac{ 6 x^2y(3x^3y^2+y^3+4 x^4) }{ 3xy }\] =?

OpenStudy (anonymous):

SO, 18x^5y^3 / 3xy 6x^2y^4/3xy 24x^6y = 6x^4y^2+2xy^3+8x^5 ?